# Proportionality for Constrained Public Decisions

### v1

@@ 4 Justified Representation with Cohesiveness
First, consider $|T|=1$. Since $|S|\leqslant|T|$ and $S\neq\emptyset$, we have $|S|=1$ for any $T$-cohesive group (which is thus of size $|N^{\prime}|\geqslant\nicefrac{{n}}{{2}}$) wishing to perform a $(S,\boldsymbol{w})$-deviation from $\boldsymbol{w}$ to some other feasible outcome $\boldsymbol{w}^{\prime}\in\mathcal{C}$. If there is a voter $i\in N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant 1$, group $N^{\prime}$ would be sufficiently satisfied—therefore, $\text{cohEJR}_{\mathcal{C}}$ is ensured and we are done. Otherwise, we have that all $i\in N^{\prime}$ are unanimous and that $u_{i}(\boldsymbol{w})=0$; hence, $\boldsymbol{b}_{i}=(1-d_{x},1-d_{y})$ for all $i\in N^{\prime}$. There are two possible outcomes (deviations) that differ from $\boldsymbol{w}$ in only one coordinate. If neither outcome is in $\mathcal{C}$, then no feasible deviation is possible for $N^{\prime}$ and we are done. Otherwise, assume without loss of generality that $\boldsymbol{w}^{\prime}=(1-d_{x},d_{y})\in\mathcal{C}$. Now, if there is a voter $i\in N\setminus N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant 1$, then we are done (as the group $N\setminus N^{\prime}$ would be sufficiently satisfied if it were $T$-cohesive for $|T|=1$). Else, it means that all voters $j\in N\setminus N^{\prime}$ are unanimous on ballot $\boldsymbol{b}_{j}=(d_{x},1-d_{y})$. But then, since $\mathcal{C}$ satisfies property NFD, there exists some outcome $\boldsymbol{w}^{\prime\prime}\in\mathcal{C}$ such that $\boldsymbol{w}^{\prime\prime}_{2}=1-d_{y}$. Then, $u_{i}(\boldsymbol{w}^{\prime\prime})\geqslant 1$ for all $i\in N$ and no deviation is possible.
@@
Finally, consider $|T|=m=2$. In order for a group $N^{\prime}$ that is $T$-cohesive to have a ($S,\boldsymbol{w}$)-deviation for $|S|\leqslant|T|$, it must be the case that $N^{\prime}=N$, and $u_{i}(\boldsymbol{w})=0$ for all $i\in N$. By property NFD, there must be some outcome $\boldsymbol{w}^{\prime}\neq\boldsymbol{w}\in\mathcal{C}$, and thus $u_{i}(\boldsymbol{w}^{\prime})\geqslant 1$ for all $i\in N$.
@@
For $|T|=1$, suppose by contradiction that for all $\boldsymbol{w}\in\mathcal{C}$, there is some voter group $N^{\prime}$ such that $|N^{\prime}|\geqslant\nicefrac{{n}}{{3}}$ and each voter in $N^{\prime}$ has satisfaction of 00. Thus, for all voters $i\in N^{\prime}$ we have $\boldsymbol{b}_{i}=\boldsymbol{\bar{w}}$. Moreover, for a $T$-cohesive group with an ($S,\boldsymbol{w}$)-deviation for $|S|=|T|=1$ to be possible, there has to exist a $\boldsymbol{w}^{\prime}\in\mathcal{C}$ whose decisions differ from $\boldsymbol{w}$ in exactly one issue, i.e., $\text{\rm Agr}(\boldsymbol{w},\boldsymbol{w}^{\prime})=2$. To fit all these disjoint $T$-cohesive groups for $|T|=1$, one for each outcome in $\mathcal{C}$, it must be that $n\geqslant|\mathcal{C}|\cdot\nicefrac{{n}}{{3}}$, hence $|\mathcal{C}|\leqslant 3$ must hold. If $|\mathcal{C}|=1$, the NFD property cannot be met. If $|\mathcal{C}|=2$, then the two feasible outcomes cannot differ in the decision of only one issue while also satisfying the NFD property. For $|\mathcal{C}|=3$, to get a $T$-cohesive voter group with an ($S,\boldsymbol{w}$)-deviation for $|S|=|T|=1$ at every $\boldsymbol{w}\in\mathcal{C}$, the three feasible outcomes must differ by at most one decision, contradicting the NFD property.
@@
For $|T|=2$, we only consider ($S,\boldsymbol{w}$)-deviations from a $T$-cohesive group $N^{\prime}$ with $|S|\in\{1,2\}$. Consider the case of $|S|=1$. W.l.o.g., assume $\boldsymbol{w}=(0,0,0)$ and assume that there exists a $T$-cohesive group $N^{\prime}$ (where $|N^{\prime}|\geqslant n\cdot\nicefrac{{2}}{{3}}$) with every voter having satisfaction $<2$, with an ($S,\boldsymbol{w}$)-deviation towards outcome, e.g., $\boldsymbol{w}^{\prime}=(1,0,0)$. The case for $(0,1,0)$ is similar. Now one of the $N^{\prime}$ voters has satisfaction of $2$. If $\nicefrac{{n}}{{3}}$ voters now have an allowable deviation (satisfaction of 00 with the current outcome), by NFD one of the outcomes $\{010,111,011,110\}$ must be in $\mathcal{C}$. Observe that any of them provides satisfaction of at least $2$ to all cohesive groups of $|T|=2$, and at least satisfaction $1$ to every cohesive group of $|T|=1$. Now we look at the case for $|S|=2$. W.l.o.g., consider the outcome $\boldsymbol{w}=(0,0,0)$ and assume that there exists a $T$-cohesive group $N^{\prime}$ (where $|N^{\prime}|\geqslant n\cdot\nicefrac{{2}}{{3}}$) with an ($S,\boldsymbol{w}$)-deviation towards outcome, e.g., $\boldsymbol{w}^{\prime}=(1,1,0)$. Thus, there is some voter $i$ in $N^{\prime}$ with satisfaction $u_{i}(\boldsymbol{w}^{\prime})\geqslant 2$. At this point, the only possible further ($S,\boldsymbol{w}$)-deviation could arise for $|S|=1$ in case there are $\nicefrac{{n}}{{3}}$ voters in $N\setminus{N^{\prime}}$ each have a satisfaction of 0 for $\boldsymbol{w}^{\prime}$, i.e., each has the ballot $(0,0,1)$ and either one of the outcomes in $\{(1,0,0),(0,1,0),(1,1,1)\}$ is in $\mathcal{C}$. Now take instead that $u_{i}(\boldsymbol{w}^{\prime})=2$ and consider two cases where either voter $i$ agrees or disagrees with the voters in $N\setminus{N^{\prime}}$ on the decision of issue $a_{3}$. First, assume that voter $i\in N^{\prime}$ agrees with the voters in $N\setminus{N^{\prime}}$ on issue $a_{3}$ (so voter $i$ had the ballot $\boldsymbol{b}_{i}=(1,1,1)$). Then if either $(0,1,1)\in\mathcal{C}$ or $(1,1,1)\in\mathcal{C}$ holds, we have that $\text{cohEJR}_{\mathcal{C}}$ is provided. And if $(0,0,1)\in\mathcal{C}$ holds, then voters in $N\setminus{N^{\prime}}$ are entirely satisfied and the voters in $N^{\prime}$ may only have an ($S,\boldsymbol{w}$)-deviation for $|S|\leqslant|T|=2$ if either $(0,1,1)\in\mathcal{C}$ or $(1,1,1)\in\mathcal{C}$ holds (as they only ‘flip’ issues they disagree with), which means that $\text{cohEJR}_{\mathcal{C}}$ is provided. In the second case, assume that voter $i\in N^{\prime}$ disagrees with the voters in $N\setminus{N^{\prime}}$ on issue $a_{3}$ and so, voter $i$ had the ballot $\boldsymbol{b}_{i}=(1,1,0)$. This means that $u_{i}(\boldsymbol{w}^{\prime})=3$ holds, hence, any outcome that the voters in $N\setminus{N^{\prime}}$ propose given they have an ($S,\boldsymbol{w}$)-deviation for $|S|=1$, would be one that provides $\text{cohEJR}_{\mathcal{C}}$.
@@
We leave it open whether the above result holds if we do not restrict our view to binary election instances. Unfortunately, the good news ends there as we provide a example showing that $\text{cohJR}_{\mathcal{C}}$ cannot be guaranteed when do not have $m\leqslant 3$ (even for binary election instances).
@@
There exists an election instance $(\boldsymbol{B},\mathcal{C})$ where $m>3$ and the constraint $\mathcal{C}$ has the NFD property but no $\text{cohJR}_{\mathcal{C}}$ outcome exists.
@@
Suppose there is a binary election instance with a constraint $\mathcal{C}=\{\boldsymbol{w}_{1},\boldsymbol{w}_{2},\boldsymbol{w}_{3},\boldsymbol{w}_{4}\}$ for $m=8$ such that $\boldsymbol{w}_{1}=(0,0,0,\ldots,0)$, $\boldsymbol{w}_{2}=(0,0,1,\ldots,1)$, $\boldsymbol{w}_{3}=(1,1,1,\ldots,1)$, $\boldsymbol{w}_{4}=(1,1,0,\ldots,0)$. Consider now a profile of four voters where $\boldsymbol{b}_{i}=\boldsymbol{w}_{i}$. Given that $m=8$, note that for every outcome $\boldsymbol{w}\in\mathcal{C}$, there exists some voter that deserves $2$ in satisfaction by being $T$-cohesive for $|T|=2$ with an $(S,\boldsymbol{w})$-deviation but with zero in satisfaction. And by $\text{cohJR}_{\mathcal{C}}$, such a voter would be entitled to at least $1$ in satisfaction, so there is no outcome in $\mathcal{C}$ that provides $\text{cohJR}_{\mathcal{C}}$. ∎
We now turn our attention towards a weakening of $\text{cohEJR}_{\mathcal{C}}$ that takes inspiration from EJR-1 studied in the context of participatory budgeting.
@@
Given an election $(\boldsymbol{B},\mathcal{C})$, an outcome $\boldsymbol{w}$ provides $\text{cohEJR}_{\mathcal{C}}$-1 if for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w})$-deviation for some $S\subseteq T$, there exists a voter $i\in N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant|T|-1$.
As $\text{cohEJR}_{\mathcal{C}}$ implies $\text{cohEJR}_{\mathcal{C}}$-1, the results of Propositions 4 and 5 immediately apply to $\text{cohEJR}_{\mathcal{C}}$-1.
@@
Note that for the computational result for $\text{cohEJR}_{\mathcal{C}}$ in Proposition 2, a simple alteration of the proof given for Proposition 2 (replacing the value $|T|$ with $|T|-1$ in the final satisfaction check) yields a corresponding computational result for $\text{cohEJR}_{\mathcal{C}}$-1.
@@
Given an election instance $(\boldsymbol{B},\mathcal{C})$ and an outcome $\boldsymbol{w}\in\mathcal{C}$, there exists an algorithm that decides in $O((\max_{t\in[m]}|D_{t}|)^{m}|\mathcal{C}|^{3}mn)$ time whether outcome $\boldsymbol{w}$ provides $\text{cohEJR}_{\mathcal{C}}$-1.
For the result of stating that $\text{cohEJR}_{\mathcal{C}}$ can be provided when $m=2$ given that NFD holds (see Proposition 5), we can show something stronger for $\text{cohEJR}_{\mathcal{C}}$-1 by dropping the assumption that the NFD property holds.
@@
Consider an election over two issues, where a $T$-cohesive group of voters has an $(S,\boldsymbol{w})$-deviation for some outcome $\boldsymbol{w}$, as per Definition 6. Observe that, when $m=2$, $(S,\boldsymbol{w})$-deviation are only possible for $|S|\in\{1,2\}$. Take a $T$-cohesive group $N^{\prime}$ for $|T|=1$ with an $(S,\boldsymbol{w})$-deviation from $\boldsymbol{w}$ to some other feasible outcome $\boldsymbol{w}^{\prime}\in\mathcal{C}$. Even if $u_{i}(\boldsymbol{w})=0$ for every voter $i\in N^{\prime}$, we have $u_{i}(\boldsymbol{w})\geqslant|T|-1=1-1=0$, and thus $\text{cohEJR}_{\mathcal{C}}$-1 is satisfied. Take now a $T$-cohesive group $N^{\prime}$ for $|T|=2$: for them to deviate, it must be the case that $N^{\prime}=N$, and $u_{i}(\boldsymbol{w})=0$ for all $i\in N$. If they have an $(S,\boldsymbol{w})$-deviation for $|S|=|T|=2$, the outcome $\boldsymbol{w^{\prime}}$ they wish to deviate to must increase the satisfaction of each voter by at least $1$, which thus satisfies $u_{i}(\boldsymbol{w})\geqslant|T|-1=2-1=1$, and thus $\text{cohEJR}_{\mathcal{C}}$-1. ∎
@@
There exists an election instance $(\boldsymbol{B},\mathcal{C})$ where $m>3$ and the constraint $\mathcal{C}$ has the NFD property but there exists no outcome that provides $\text{cohEJR}_{\mathcal{C}}$-1.
We demonstrate that the challenge of satisfying $\text{cohEJR}_{\mathcal{C}}$-1 lies in the constraints. To do so, we show that in the setting without constraints, it is always possible to find an outcome that provides $\text{cohEJR}_{\mathcal{C}}$-1. To do so, we define the constrained version of MES that has been studied for the public-decision setting without constraints. Our adaptation allows for the prices associated with fixing the outcome’s decisions on issues to vary. This contrasts with the unconstrained MES that fixes the prices of every issue’s decision to $n$ from the onset. And this pricing is determined by a particular pricing type $\lambda$.
@@
The rule runs for at most $m$ rounds. Each voter has a budget of $m$. In every round, for every undecided issue $a_{t}$ in a partial outcome $\boldsymbol{w}^{*}$, we identify those issue-decision pairs $(a_{t},d)$ where fixing some decision $d\in D_{t}$ on issue $a_{t}$ allows for a feasible outcome to be returned in future rounds. If no such issue-decision pair exists, then the rule stops. Otherwise, for every such pair $(a_{t},d)$, we calculate the minimum value for $\rho_{(a_{t},d)}$ such that if each voter in $N(a_{t},d)$ were to pay either $\rho_{(a_{t},d)}$ or the remainder of their budget, then these voters could afford to pay the price $\lambda(a_{t},d)$ (determined by the pricing type $\lambda$). If there exists no such value for $\rho_{(a_{t},d)}$, then we say that the issue-decision pair $(a_{t},d)$ is not affordable in round, and if in a round, there are no affordable issue-decision pairs, the rule stops. Otherwise, we update $\boldsymbol{w}^{*}$ by setting decision $d$ on issue $a_{t}$ for the pair $(a_{t},d)$ with a minimal value $\rho_{(a_{t},d)}$ (breaking ties arbitrarily, if necessary) and have each voter in $N(a_{t},d)$ either paying $\rho_{(a_{t},d)}$, or the rest of their budget. Note that $\text{MES}_{\mathcal{C}}$ may terminate with not all issues being decided and we assume that all undecided issues are decided arbitrarily.

### v2

@@ 4 Justified Representation with Cohesiveness
First, consider $|T|=1$. Since $|S|\leqslant|T|$ and $S\neq\emptyset$, we have $|S|=1$ for any $T$-cohesive group (which is thus of size $|N^{\prime}|\geqslant\nicefrac{{n}}{{2}}$) wishing to perform an $(S,\boldsymbol{w},\mathcal{C})$-deviation from $\boldsymbol{w}$ to some other feasible outcome $\boldsymbol{w}^{\prime}\in\mathcal{C}$. If there is a voter $i\in N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant 1$, group $N^{\prime}$ would be sufficiently satisfied and $\text{cohEJR}_{\mathcal{C}}$ would be ensured. Otherwise, we have that $u_{i}(\boldsymbol{w})=0$ for all $i\in N^{\prime}$ and they are unanimous on both issues, i.e., $\boldsymbol{b}_{i}=(1-d_{1},1-d_{2})$ for all $i\in N^{\prime}$. There are just two possible outcomes that differ from $\boldsymbol{w}$ on only one issue. If neither outcome is in $\mathcal{C}$, then no feasible deviation is possible for $N^{\prime}$ and we are done. Otherwise, assume without loss of generality that $\boldsymbol{w}^{\prime}=(1-d_{1},d_{2})\in\mathcal{C}$. Now, if there is a voter $i\in N\setminus N^{\prime}$ such that $u_{i}(\boldsymbol{w}^{\prime})\geqslant 1$, then we are done (as the group $N\setminus N^{\prime}$ would be sufficiently satisfied if it were $T$-cohesive for $|T|=1$). Else, it means that all voters $j\in N\setminus N^{\prime}$ are unanimous on ballot $\boldsymbol{b}_{j}=(d_{1},1-d_{2})$. But then, since $\mathcal{C}$ satisfies property NFD, there exists some outcome $\boldsymbol{w}^{\prime\prime}\in\mathcal{C}$ such that $\boldsymbol{w}^{\prime\prime}_{2}=1-d_{2}$. Then, $u_{i}(\boldsymbol{w}^{\prime\prime})\geqslant 1$ for all $i\in N$ and no deviation is possible.
@@
Now, consider $|T|=m=2$. For a voter group $N^{\prime}$ to be $T$-cohesive, they must unanimously agree on both issues and their size must be $|N^{\prime}|=n$. In order for the group $N^{\prime}$ to have an ($S,\boldsymbol{w},\mathcal{C}$)-deviation for $|S|\leqslant|T|$, it must be the case that $u_{i}(\boldsymbol{w})=0$ for all $i\in N^{\prime}$. By property NFD, there must be some outcome $\boldsymbol{w}^{\prime}\neq\boldsymbol{w}\in\mathcal{C}$, and thus $u_{i}(\boldsymbol{w}^{\prime})\geqslant 1$ for all $i\in N$.
@@
For $|T|=1$, suppose by contradiction that for all $\boldsymbol{w}\in\mathcal{C}$, there is some voter group $N^{\prime}$ such that $|N^{\prime}|\geqslant\nicefrac{{n}}{{3}}$ and each voter in $N^{\prime}$ has satisfaction of 0. Thus, for all voters $i\in N^{\prime}$ we have $\boldsymbol{b}_{i}=\boldsymbol{\bar{w}}$. Moreover, for a $T$-cohesive group with an ($S,\boldsymbol{w}$)-deviation for $|S|=|T|=1$ to be possible, there has to exist a $\boldsymbol{w}^{\prime}\in\mathcal{C}$ whose decisions differ from $\boldsymbol{w}$ in exactly one issue, i.e., $\text{\rm Agr}(\boldsymbol{w},\boldsymbol{w}^{\prime})=2$. To fit all these disjoint $T$-cohesive groups for $|T|=1$, one for each outcome in $\mathcal{C}$, it must be that $n\geqslant|\mathcal{C}|\cdot\nicefrac{{n}}{{3}}$, hence $|\mathcal{C}|\leqslant 3$ must hold. If $|\mathcal{C}|=1$, the NFD property cannot be met. If $|\mathcal{C}|=2$, then the two feasible outcomes cannot differ in the decision of only one issue while also satisfying the NFD property. For $|\mathcal{C}|=3$, to get a $T$-cohesive voter group with an ($S,\boldsymbol{w}$)-deviation for $|S|=|T|=1$ at every $\boldsymbol{w}\in\mathcal{C}$, the three feasible outcomes must differ by at most one decision, contradicting the NFD property.
@@
For $|T|=2$, we only consider ($S,\boldsymbol{w}$)-deviations from a $T$-cohesive group $N^{\prime}$ with $|S|\in\{1,2\}$. Consider the case of $|S|=1$. W.l.o.g., assume that $\boldsymbol{w}=(0,0,0)$ and that there exists a $T$-cohesive group $N^{\prime}$ (where $|N^{\prime}|\geqslant n\cdot\nicefrac{{2}}{{3}}$) with every voter having satisfaction $u_{i}(\boldsymbol{w})<2$, with an ($S,\boldsymbol{w}$)-deviation towards some outcome, e.g., $\boldsymbol{w}^{\prime}=(1,0,0)$. So $u_{i}(\boldsymbol{w}^{\prime})\geqslant 1$ holds for at least one voter $i\in N^{\prime}$. If $N^{\prime}$ has no ($S,\boldsymbol{w}^{\prime}$)-deviation, then there are not a witness to a violation of $\text{cohEJR}_{\mathcal{C}}$. Otherwise, suppose no voter in $N^{\prime}$ has satisfaction of $2$ (so all voters in $N^{\prime}$ have a ballot $(1,1,1)$) and the group $N^{\prime}$ has an ($S,\boldsymbol{w}^{\prime}$)-deviation, to outcome $\boldsymbol{w}^{\prime\prime}=(1,1,0)$. Then all voters in $N^{\prime}$ have satisfaction of at least $2$ with outcome $\boldsymbol{w}^{\prime\prime}$. Suppose that the remaining $\nicefrac{{n}}{{3}}$ of the voters $N\setminus N^{\prime}$ have an ($S,\boldsymbol{w}^{\prime\prime}$)-deviation (so each of these voters derives zero satisfaction from outcome $\boldsymbol{w}^{\prime\prime}$ and all disagree with group $N^{\prime}$ on the first two issues $a_{1}$ and $a_{2}$). This must be a deviation of size $|S|=1$ for the $\nicefrac{{n}}{{3}}$ of the voters $N\setminus N^{\prime}$ to demand it. This deviation can only be to outcome $(1,1,1)$, $(0,1,0)$ or $(1,0,0)$. By the NFD property, we know that one of the outcomes $\{(0,0,1),(0,1,1),(1,0,1),(1,1,1)\}$ must be in $\mathcal{C}$. See that for each of the outcomes $(1,1,1)$, $(1,0,1)$ and $(0,1,1)$, some voter in $N^{\prime}$ gets at least satisfaction of $2$ while all voters in $N\setminus N^{\prime}$ get at least $1$ in satisfaction, thus, $\text{cohEJR}_{\mathcal{C}}$ is satisfied. Now, out of these outcomes $\{(0,0,1),(0,1,1),(1,0,1),(1,1,1)\}$, if only outcome $\boldsymbol{w}^{\prime\prime\prime}=(0,0,1)$ is in $\mathcal{C}$, then group $N^{\prime}$ has no ($S,\boldsymbol{w}^{\prime\prime\prime}$)-deviations of size $|S|=1$ as this is only possible to one of $(0,1,1)$ or $(1,0,1)$. So in this case too, $\text{cohEJR}_{\mathcal{C}}$ is satisfied.
Now we look at the case for $|S|=2$. W.l.o.g., consider the outcome $\boldsymbol{w}=(0,0,0)$ and assume that there exists a $T$-cohesive group $N^{\prime}$ (where $|N^{\prime}|\geqslant n\cdot\nicefrac{{2}}{{3}}$) with an ($S,\boldsymbol{w}$)-deviation towards outcome, e.g., $\boldsymbol{w}^{\prime}=(1,1,0)$. Thus, there is some voter $i$ in $N^{\prime}$ with satisfaction $u_{i}(\boldsymbol{w}^{\prime})\geqslant 2$. At this point, the only possible further ($S,\boldsymbol{w}$)-deviation could arise for $|S|=1$ in case there are $\nicefrac{{n}}{{3}}$ voters in $N\setminus{N^{\prime}}$ each have a satisfaction of 0 for $\boldsymbol{w}^{\prime}$, i.e., each has the ballot $(0,0,1)$ and either one of the outcomes in $\{(1,0,0),(0,1,0),(1,1,1)\}$ is in $\mathcal{C}$. Now take instead that $u_{i}(\boldsymbol{w}^{\prime})=2$ and consider two cases where either voter $i$ agrees or disagrees with the voters in $N\setminus{N^{\prime}}$ on the decision of issue $a_{3}$. First, assume that voter $i\in N^{\prime}$ agrees with the voters in $N\setminus{N^{\prime}}$ on issue $a_{3}$ (so voter $i$ had the ballot $\boldsymbol{b}_{i}=(1,1,1)$). Then if either $(0,1,1)\in\mathcal{C}$ or $(1,1,1)\in\mathcal{C}$ holds, we have that $\text{cohEJR}_{\mathcal{C}}$ is provided. And if $(0,0,1)\in\mathcal{C}$ holds, then voters in $N\setminus{N^{\prime}}$ are entirely satisfied and the voters in $N^{\prime}$ may only have an ($S,\boldsymbol{w}$)-deviation for $|S|\leqslant|T|=2$ if either $(0,1,1)\in\mathcal{C}$ or $(1,1,1)\in\mathcal{C}$ holds (as they only ‘flip’ issues they disagree with), which means that $\text{cohEJR}_{\mathcal{C}}$ is provided. In the second case, assume that voter $i\in N^{\prime}$ disagrees with the voters in $N\setminus{N^{\prime}}$ on issue $a_{3}$ and so, voter $i$ had the ballot $\boldsymbol{b}_{i}=(1,1,0)$. This means that $u_{i}(\boldsymbol{w}^{\prime})=3$ holds, hence, any outcome that the voters in $N\setminus{N^{\prime}}$ propose given they have an ($S,\boldsymbol{w}$)-deviation for $|S|=1$, would be one that provides $\text{cohEJR}_{\mathcal{C}}$.
@@
We leave it open whether the above result holds if we do not restrict our view to binary election instances. Unfortunately, we provide an example showing that $\text{cohJR}_{\mathcal{C}}$ cannot be guaranteed in general even when the NFD property holds for binary election instances.
@@
There exists an election instance $(\boldsymbol{B},\mathcal{C})$ where $m=8$ and the constraint $\mathcal{C}$ has the NFD property but no $\text{cohJR}_{\mathcal{C}}$ outcome exists.
@@
Suppose there is a binary election instance with a constraint $\mathcal{C}=\{\boldsymbol{w}_{1},\boldsymbol{w}_{2},\boldsymbol{w}_{3},\boldsymbol{w}_{4}\}$ for $m=8$ such that $\boldsymbol{w}_{1}=(0,0,0,\ldots,0)$, $\boldsymbol{w}_{2}=(0,0,1,\ldots,1)$, $\boldsymbol{w}_{3}=(1,1,1,\ldots,1)$ and $\boldsymbol{w}_{4}=(1,1,0,\ldots,0)$. Consider now a profile of four voters where $\boldsymbol{b}_{i}=\boldsymbol{w}_{i}$. Given that $m=8$, note that for every outcome $\boldsymbol{w}\in\mathcal{C}$, there exists some voter that deserves $2$ in satisfaction by being $T$-cohesive for $|T|=2$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation but with zero in satisfaction. And by $\text{cohJR}_{\mathcal{C}}$, such a voter would be entitled to at least $1$ in satisfaction, so there is no outcome in $\mathcal{C}$ that provides $\text{cohJR}_{\mathcal{C}}$. ∎
We now turn our attention towards a weakening of $\text{cohEJR}_{\mathcal{C}}$ that takes inspiration from the ‘up-to-one’ relaxation of EJR studied in the context of participatory budgeting (Peters et al., 2021b; Rey and Maly, 2023).
@@
Given an election $(\boldsymbol{B},\mathcal{C})$, an outcome $\boldsymbol{w}$ provides $\text{cohEJR}_{\mathcal{C}}$-1 if for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some non-empty $S\subseteq T$, there exists a voter $i\in N^{\prime}$ such that:
$u_{i}(\boldsymbol{w})\geqslant|T|-1$
Intuitively, this axiom states that a fair outcome would be one where all groups that are can demand a certain level of satisfaction as defined by $\text{cohEJR}_{\mathcal{C}}$, are a single decision away from obtaining their deserved satisfaction. As $\text{cohEJR}_{\mathcal{C}}$ implies $\text{cohEJR}_{\mathcal{C}}$-1, the results of Propositions 4 and 5 immediately apply to $\text{cohEJR}_{\mathcal{C}}$-1.
@@
Note that for the computational result for $\text{cohEJR}_{\mathcal{C}}$ in Proposition 2, a simple alteration of the proof given for Proposition 2 (replacing the value $|T|$ with |T|1|T|-1 in the final satisfaction check) yields a corresponding computational result for $\text{cohEJR}_{\mathcal{C}}$-1 that is similarly (negatively) impacted by the number of issues in the election instance.
@@
Given an election instance $(\boldsymbol{B},\mathcal{C})$ and an outcome $\boldsymbol{w}\in\mathcal{C}$, there exists an algorithm that decides in $O(2(\max_{t\in[m]}|D_{t}|)^{m}|\mathcal{C}|^{3}mn)$ time whether outcome $\boldsymbol{w}$ provides $\text{cohEJR}_{\mathcal{C}}$-1.
Next, for the result that shows that $\text{cohEJR}_{\mathcal{C}}$ can be provided when $m=2$ given that NFD holds (see Proposition 5), we can show something stronger for $\text{cohEJR}_{\mathcal{C}}$-1 by dropping the assumption that the NFD property holds.
@@
Consider an election over two issues, where a $T$-cohesive group of voters has an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some outcome $\boldsymbol{w}$, as per Definition 6. Observe that, when $m=2$, $(S,\boldsymbol{w},\mathcal{C})$-deviation are only possible for $|S|\in\{1,2\}$. Take a $T$-cohesive group $N^{\prime}$ for $|T|=1$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation from $\boldsymbol{w}$ to some other feasible outcome $\boldsymbol{w}^{\prime}\in\mathcal{C}$. Even if $u_{i}(\boldsymbol{w})=0$ for every voter $i\in N^{\prime}$, we have $u_{i}(\boldsymbol{w})\geqslant|T|-1=1-1=0$, and thus $\text{cohEJR}_{\mathcal{C}}$-1 is satisfied. Take now a $T$-cohesive group $N^{\prime}$ for $|T|=2$: for them to deviate, it must be the case that $N^{\prime}=N$, and $u_{i}(\boldsymbol{w})=0$ for all $i\in N$. If they have an $(S,\boldsymbol{w},\mathcal{C})$-deviation for $|S|=|T|=2$, the outcome $\boldsymbol{w^{\prime}}$ they wish to deviate to must increase the satisfaction of each voter by at least $1$, which thus satisfies $u_{i}(\boldsymbol{w})\geqslant|T|-1=2-1=1$, and thus $\text{cohEJR}_{\mathcal{C}}$-1. ∎
@@
There exists an election instance $(\boldsymbol{B},\mathcal{C})$ where $m=8$ and the constraint $\mathcal{C}$ has the NFD property but there exists no outcome that provides $\text{cohEJR}_{\mathcal{C}}$-1.
We demonstrate that the challenge of satisfying $\text{cohEJR}_{\mathcal{C}}$-1 lies in the constraints. In fact, we show that in the setting without constraints, it is always possible to find an outcome that provides $\text{cohEJR}_{\mathcal{C}}$-1. To do so, we define the constrained version of MES that has been studied for the public-decision setting without constraints. Our adaptation allows for the prices associated with fixing the outcome’s decisions on issues to vary. This contrasts with the unconstrained MES that fixes the prices of every issue’s decision to $n$ from the onset. This pricing is determined by a particular pricing type $\lambda:\mathcal{I}\times(D_{1}\cup\ldots\cup D_{m})\rightarrow\mathbb{R}_{\geqslant 0}$ which maps an issue-decision pair to a non-negative price.
@@
The rule runs for at most $m$ rounds. Each voter has a budget of $m$. In every round, for every undecided issue $a_{t}$ in a partial outcome $\boldsymbol{w}^{*}$, we identify those issue-decision pairs $(a_{t},d)$ where fixing some decision $d\in D_{t}$ on issue $a_{t}$ allows for a feasible outcome to be returned in future rounds. Otherwise, for every such pair $(a_{t},d)$, we calculate the minimum value for $\rho_{(a_{t},d)}$ such that if each voter in $N(a_{t},d)$ were to pay either $\rho_{(a_{t},d)}$ or the remainder of their budget, then these voters could afford to pay the price $\lambda(a_{t},d)$ (determined by the pricing type $\lambda$). If there exists no such value for $\rho_{(a_{t},d)}$, then we say that the issue-decision pair $(a_{t},d)$ is not affordable in round, and if in a round, there are no affordable issue-decision pairs, the rule stops. Otherwise, we update $\boldsymbol{w}^{*}$ by setting decision $d$ on issue $a_{t}$ for the pair $(a_{t},d)$ with a minimal value $\rho_{(a_{t},d)}$ (breaking ties arbitrarily, if necessary) and have each voter in $N(a_{t},d)$ either paying $\rho_{(a_{t},d)}$, or the rest of their budget. Note that $\text{MES}_{\mathcal{C}}$ may terminate with not all issues being decided and we assume that all undecided issues are decided arbitrarily.